How do you solve 3/4 • 5/6
Found 2 solutions by ptth176, Alan3354:
Answer by ptth176(50)

3/4 • 5/6=(3x5)/(4.6)=15/24

Answer by Alan3354(69443)

How do you solve 3/4 • 5/6
--------------
You don't solve it, there's no variable, such as x, to solve for.
Equations are solved.
You can multiply the 2 numbers
----------
3/4 • 5/6 = 15/24
= 5/8